# 3D-free meets 3D priors: Novel View Synthesis from a Single Image with Pretrained Diffusion Guidance

### v4

@@ Abstract
Recent 3D novel view synthesis (NVS) methods often require extensive 3D data for training, and also typically lack generalization beyond the training distribution. Moreover, they tend to be object centric and struggle with complex and intricate scenes. Conversely, 3D-free methods can generate text-controlled views of complex, in-the-wild scenes using a pretrained stable diffusion model without the need for a large amount of 3D-based training data, but lack camera control. In this paper, we introduce a method capable of generating camera-controlled viewpoints from a single input image, by combining the benefits of 3D-free and 3D-based approaches. Our method excels in handling complex and diverse scenes without extensive training or additional 3D and multiview data. It leverages widely available pretrained NVS models for weak guidance, integrating this knowledge into a 3D-free view synthesis style approach, along with enriching the CLIP vision-language space with 3D camera angle information, to achieve the desired results. Experimental results demonstrate that our method outperforms existing models in both qualitative and quantitative evaluations, achieving high-fidelity, consistent novel view synthesis at desired camera angles across a wide variety of scenes while maintaining accurate, natural detail representation and image clarity across various viewpoints. We also support our method with a comprehensive analysis of 2D image generation models and the 3D space, providing a solid foundation and rationale for our solution.
Figure 1: Our model is capable of generating high quality camera-controlled images at specific azimuth and elevation angles for a variety of complex scenes, all without requiring extra 3D datasets or extensive training. The image in the bottom right corner showcases the output from the 3D-based baseline, Zero123++ Shi et al. (2023a), created from a designated angle.
@@
1 Introduction
Novel-view synthesis plays a pivotal role in numerous real-world applications, including 3D environments, augmented reality (AR), virtual reality (VR), and autonomous driving. Recent advancements in diffusion-based methods, such as Zero-1-to-3 (Zero123) Liu et al. (2023a) and Zero123++ Shi et al. (2023a), along with NeRFs Tancik et al. (2022); Zhou & Tulsiani (2023); Deng et al. (2023) Gaussian Splatting Li et al. (2024); Zhu et al. (2023), and so on Sargent et al. (2023); Tung et al. (2025); Van Hoorick et al. (2025); Burgess et al. (2023); Wiles et al. (2020); Shen et al. (2021); Tucker & Snavely (2020). They have significantly propelled the field forward. Some techniques enable the specification of camera angles and the sampling of novel-view images from precise viewpoints. However, diffusion-based methods remain largely object-centric and may struggle to generalize to complex scenes with intricate backgrounds. They also require extensive 3D object datasets for training. In contrast, NeRF and Gaussian Splatting methods can handle complex scenes but depend on multi-view information to construct 3D models. Therefore, achieving novel-view synthesis from a single image in a data-efficient manner, without relying on additional 3D, multi-view, or depth information, is highly advantageous.
On the other hand, 3D-free methods such as DreamBooth and other recent models Kothandaraman et al. (2023b; a); Ruiz et al. (2023) aim to intelligently extract the rich 3D knowledge embedded in text-to-2D image diffusion models, such as stable diffusion Rombach et al. (2022); Podell et al. (2023), to generate text-controlled views from complex, real-world input images without needing additional multi-view or 3D information for fine-tuning or inference. Among these, HawkI stands out as the current best 3D-free model, demonstrating superior capability in utilizing embedded 3D knowledge for high-quality, text-controlled image synthesis. However, despite its excellence, HawkI, like other 3D-free methods, lacks the ability to precisely control camera angles when generating novel-view images. Ideally, we aim for both data-efficient novel view synthesis and camera controllability, which is the primary focus of this paper.
@@ 1 Introduction
Using these insights, along with the established knowledge that 3D-based methods such as HawkI enable precise camera control and 3D-free methods like Zero123++ offer generalizability and data efficiency, we propose a simple approach for novel-view synthesis that generates camera-controlled novel views at specified azimuth and elevation angles from a single input image, without requiring 3D datasets or extensive training. Our method utilizes information from off-the-shelf pretrained model, specifically using Zero123++ Shi et al. (2023a), a plug-and-play model, in conjunction with the pretrained stable diffusion model. The process employs a 3D-free HawkI-style optimization procedure during inference to achieve the desired outcomes, utilizing information from 3D-based methods as pseudo or weak guidance images. To improve viewpoint consistency—an area where the CLIP model lacks information—we introduce a regularization loss term. This term promotes alignment between the target angle embedding (which captures elevation and azimuth data) and the optimized embedding. By integrating 3D angular information within the CLIP space through the 3D prior image, we reinforce the specified camera viewpoint in the generated images. We validate our approach through extensive qualitative and quantitative comparisons across various metrics that assess text consistency and fidelity w.r.t. input image.
@@
In summary, the contributions of this paper are as follows: (1) We present a novel approach for novel view synthesis that allows for precise camera control, especially effective for complex images with multiple objects and detailed backgrounds. Our method harnesses insights from pretrained 3D models within a 3D-free framework, removing the necessity for additional multi-view or 3D data during both training and inference, effectively combining the advantages of both approaches. (2) We provide an analysis of the CLIP model’s understanding of 3D space and the role of guidance images in 3D-free methods. This analysis supports our solution by highlighting the importance of using priors from pretrained 3D models to enhance viewpoint information in 3D space, utilizing the 3D prior image as a guiding factor for our task. (3) We present comprehensive qualitative and quantitative results on various synthetic and real images, demonstrating significant improvements over baseline 3D-based and 3D-free methods in terms of text consistency and fidelity relative to the input images. Our model’s results maintains consistent, accurate, natural detail representation and image clarity across various viewpoints. Also, our model outperforms the lowest-performing model by 0.1712 in LPIPS (HawkI-Syn $(-20^{\circ},210^{\circ})$ in Table 5), which is 5.2 times larger than the 0.033 gap of Zero123++ in comparison to the lowest-performing model (Table 1 in Shi et al. (2023a)).
2 Prior Work on 3D and 3D-free Approaches for NVS
@@
Recent research has increasingly focused on novel view synthesis using diffusion models Chen et al. (2021); Mildenhall et al. (2021); Shi et al. (2021); Gu et al. (2023). Approaches for 3D generation Chen et al. (2024); Lin et al. (2023); Poole et al. (2022); Raj et al. (2023); Xu et al. (2023); Gao et al. (2024); Park et al. (2017) often rely on text for reconstruction and require substantial multi-view and 3D data Shi et al. (2023b); Wang & Shi (2023); Yang et al. (2024); Liu et al. (2023b); Höllein et al. (2024); Jain et al. (2021); Liu et al. (2024); Shi et al. (2023c; b) for supervised learning. For instance, Zero123 Liu et al. (2023a), Magic123 Qian et al. (2023), and Zero123++ Shi et al. (2023a) utilize a pre-trained stable diffusion model Rombach et al. (2022) combined with 3D data corresponding to 800k objects to learn various camera viewpoints. In other words, they are extremely data hungry, meaning that they need extensive multi-view and 3D data to train on. Additionally, most state-of-the-art view synthesis algorithms are largely object-centric, and may not work well on complex scenes containing multiple objects or background information. This is due to the domain gap between the 3D objects data that they are typically trained on, and the inferencing image.
@@
On the other hand, Free3D Zheng & Vedaldi (2024) introduces an efficient method for synthesizing accurate 360-degree views from a single image without 3D representations. By incorporating the Ray Conditioning Normalization (RCN) layer into 2D image generators, it encodes the target view’s pose and enhances view consistency with lightweight multi-view attention layers and noise sharing. However, it still requires training on large-scale 3D datasets like Objaverse and multi-view information, and it cannot include background transformations. DreamFusion Poole et al. (2022) presents a Text-to-3D method using a NeRF and a Diffusion Model-based Text-to-2D model. It introduces a probability density distillation loss, allowing the 2D Diffusion Model to optimize image generation without needing 3D data or model modifications. DreamFusion’s key contributions are creating Text-to-3D models without 3D dataset training and utilizing a Diffusion Model. However, it cannot transform images with backgrounds or introduce elevation changes in camera-controlled images. Aerial Diffusion and HawkI Kothandaraman et al. (2023a; b) synthesize high-quality aerial view images using text and a single input image without 3D or multi-view information. They employ a pretrained text-to-2D Stable Diffusion model, achieving a balance between aerial view consistency and input image fidelity through test-time optimization and mutual information-based inference. However, Aerial Diffusion doesn’t extend well to complex scenes and has artifacts in the generated results, and HawkI struggles with controlling camera angles, detailed feature generation, and maintaining view consistency.
3 Understanding 2D models and the 3D space
@@
In this section, we use 3D-free stable diffusion based view synthesis method, HawkI Kothandaraman et al. (2023b). HawkI employs classical computer vision techniques to generate aerial view images from ground view images through a homography transformation, which acts as the guidance image for the diffusion model. We used the HawkI default setting for experiments, unless stated otherwise.
3.1 How well do CLIP models understand the 3D-space?
@@
The main reason stable-diffusion-based 2D models using 3D-free approaches struggle with camera control is their limited understanding of 3D space. While they can grasp concepts like top, bottom, and side, they lack precise camera information, such as “30 degrees to the right.” A key factor in this limitation is the CLIP model, which serves as the vision-language backbone for models like Stable Diffusion. In this section, we analyze how effectively CLIP models comprehend 3D space by examining the 3D-free stable diffusion based view synthesis method, HawkI Kothandaraman et al. (2023b). Our hypothesis is that HawkI’s capability to execute viewpoint transformations without relying on 3D data is dependent on this homography image.
@@
3.2 Importance of guidance image in 3D-free methods
@@
We conduct experiments using the images generated using the pretrained Zero123++ Shi et al. (2023a) model for guidance. In our experiments, the target text specified the desired transformation angles, but the guidance images had different angles. i.e. the guidance images introduced incorrect viewpoints. When generating an image at a $(30^{\circ},30^{\circ})$ angle, the model followed the guidance image’s suggestion, regardless of the text input. This emphasizes that the model benefits from the information in the 3D-prior model’s guidance image a lot. Our experiment highlights the importance of accurate guidance images for camera control. Figure 4 illustrates these findings.
@@
4 3D-Free Meets 3D Priors: An approach for 3D-data efficient NVS
According to the analysis, we present a method (Figure 2) for data-efficient text and camera-controlled novel-view synthesis from a single input image ($I_{input}$) and its text description ($t_{input}$) (e.g., “An ancient Egyptian pyramid in the desert,” obtained using the BLIP-2 Li et al. (2023) model). Our model eliminates the need for training data, 3D data, or multi-view data. Instead, it utilizes a pretrained text-to-2D image stable diffusion model as a strong prior, along with pretrained novel-view synthesis (NVS) models, e.g. Zero123++, for guidance. Our method combines information from the pretrained NVS model and performs a rapid inference-time optimization and inference routine to generate novel-view images of any given in-the-wild complex input scene at specified elevation ($\alpha_{elev}$) and azimuth angles ($\alpha_{azi}$). Elevation ($\alpha_{elev}$) refers to the vertical angle relative to the object, measured in degrees, and is defined based on the orientation of the input image. Similarly, azimuth angles ($\alpha_{azi}$) refer to the horizontal angle around the object, also relative to the input image. We next describe our method in detail.
4.1 Inference-time Optimization
We employ a pretrained NVS model $G$ to obtain a weak prediction, $I_{view}$, of $I_{input}$ at ($\alpha_{elev}$, $\alpha_{azi}$). This prediction is represented as $I_{view}=G(I_{input},\alpha_{elev},\alpha_{azi})$. Although $I_{view}$ is not a fully accurate depiction of the desired target, it provides weak or pseudo guidance for the model regarding the content and direction of the desired viewpoint transformation. Subsequently, we utilize the pretrained text-to-image stable diffusion model to perform inference-time optimization Kothandaraman et al. (2023b).
Across all four steps, the reconstruction loss $L$ is used to guide the optimization process, ensuring accurate reconstruction of $I_{input}$ and $I_{view}$. In Step 4, the addition of the regularization loss reinforces viewpoint consistency by aligning $e_{view}$ with $e_{target}$, thereby improving camera-controlled image generation. Details about regularization loss are mentioned in  4.1.5.
4.1.1 Step 1: Text Embedding Optimization on $I_{input}$
Initially, we enhance the CLIP embedding for $t_{input}$ with $I_{input}$ to derive $e_{optim}$ (optimized CLIP text-image embedding from $e_{input}$, which is the CLIP test embedding for $t_{input}$). This embedding is optimized to most accurately reconstruct $I_{input}$. $f$ represents the diffusion model function that maps the input latent $x_{t}$, timestep $t$, and the optimized embedding $e_{optim}$. The reconstruction is achieved by minimizing the denoising diffusion loss function $L$ Ho et al. (2020), using the frozen diffusion model UNet:
[…]
@@ 4.1.1 Step 1: Text Embedding Optimization on $I_{input}$
4.1.2 Step 2: Fine-tuning UNet on $I_{input}$
@@
4.1.3 Step 3: Text Embedding Optimization on $I_{view}$
@@
[…]
4.1.4 Step 4: Fine-tuning UNet on $I_{view}$ with Regularization Loss
@@
[…]
4.1.5 Viewpoint Regularization
@@
We camera control results by adding a regularization term between the text embedding that includes elevation and azimuth information ($e_{target}$) and the optimized text embedding ($e_{view}$) in addition to the pretrained guidance model. In addition to enriching the viewpoint knowledge of the 3D space, applying this loss is also essential to address the limitations of the 3D-prior guidance model, as the 3D-prior model does not perform very well in complex scenes. Specifically, by applying a regularization loss between the text embedding that contains angle information and the optimized text embedding, our hypothesis is that we can improve camera control results by building a model that references the guidance image as supplementary information rather than relying solely on it.
Hence, to improve viewpoint consistency, an additional regularization loss term is added to the reconstruction loss. This term introduces a constraint between the angle embedding $e_{target}$ (representing elevation and azimuth information) and the optimized embedding $e_{view}$ during this refinement. The regularization term, calculated as $L_{reg}=\|e_{view}-e_{target}\|^{2}$ encourages alignment between $e_{view}$ and the intended angle information in $e_{target}$, reinforcing the target viewpoint in the generated results.
Thus, the predicted image from the 3D-based NVS method, Zero123++, serves as weak or pseudo guidance. The optimization strategy conditions the embedding space with knowledge related to the input image and its view variants using the guidance image prior which facilitates view transformation and provides a direction for viewpoint transformation.
4.2 Inference
To generate the camera-controlled image with designated elevation ($\alpha_{elev}$) and azimuth angles ($\alpha_{azi}$), we use the target text description $t_{target}$, which varies according to the corresponding $\alpha_{elev}$ and $\alpha_{azi}$. For instance, if $\alpha_{elev}=30^{\circ}$ and $\alpha_{azi}=30^{\circ}$, $t_{target}$ can be formatted as ”View from an elevated angle of +30 degrees and an azimuth angle of +30 degrees” + $t_{input}$ (e.g., ”View from an elevated angle of +30 degrees and an azimuth angle of +30 degrees, An ancient Egyptian pyramid in the desert.”). Next, we use the finetuned diffusion model to generate the target image using $t_{target}$, along with mutual information guidance Kothandaraman et al. (2023b), which enforces similarity between the contents of the generated and input images.
5 Experiments and Results
@@
We utilize the HawkI-Syn Kothandaraman et al. (2023b) and HawkI-Real Kothandaraman et al. (2023b) datasets that feature complex scenes with multiple foreground objects and background. Both datasets provide images and text prompts to the model.
@@
We compare our method with state-of-the-art view synthesis methods: Zero123++ Shi et al. (2023a) and Stable Zero123 for 3D-based methods and HawkI Kothandaraman et al. (2023b) for 3D-free method.
@@
We employ the stable diffusion 2.1 model as the backbone for all our experiments and results. To generate the pseudo guidance images for different viewpoints, we use the pretrained Zero123++ Shi et al. (2023a) model. All images except those in HawkI-Real dataset are used at a resolution of $512\times 512$. For $I_{input}$, we train the text embedding for 1,000 iterations with the learning rate of $31e-$ 3 and the diffusion model for 500 iterations with the learning rate of $42e-$ 4. Training the text embedding for 1,000 iterations guarantees that the text embedding $e_{optim}$ is not too close to the $e_{input}$, avoiding bias towards $I_{input}$. Likewise, it is not too distant from $e_{input}$, allowing the text embedding space to capture the characteristics of $I_{input}$. Regarding $I_{view}$, we trained the text embedding for 500 iterations and the diffusion UNet for 250 iterations. We aim for $e_{view}$ to be near $e_{optim}$ and limit the diffusion model training to 250 iterations to prevent overfitting to $I_{view}$. The purpose of $I_{view}$ is to introduce variability and provide pseudo supervision rather than accurately approximating the camera control. We set the mutual information guidance hyperparameter to $61e-$ 6 and conduct inference over 50 steps.
5.1 Qualitative analysis
@@
$\{(\alpha_{elev},\alpha_{azi})\mid(30^{\circ},30^{\circ}),(-20^{\circ},210^{\circ}),(30^{\circ},270^{\circ}),\vspace*{-0.5em}(-20^{\circ},330^{\circ})\}.$

### v5

@@ Abstract
Recent 3D novel view synthesis (NVS) methods often require extensive 3D data for training, and also typically lack generalization beyond the training distribution. Moreover, they tend to be object centric and struggle with complex and intricate scenes. Conversely, 3D-free methods can generate text-controlled views of complex, in-the-wild scenes using a pretrained stable diffusion model without the need for a large amount of 3D-based training data, but lack camera control. In this paper, we introduce a method capable of generating camera-controlled viewpoints from a single input image, by combining the benefits of 3D-free and 3D-based approaches. Our method excels in handling complex and diverse scenes without extensive training or additional 3D and multiview data. It leverages widely available pretrained NVS models for weak guidance, integrating this knowledge into a 3D-free view synthesis style approach, along with enriching the CLIP vision-language space with 3D camera angle information, to achieve the desired results. Experimental results demonstrate that our method outperforms existing models in both qualitative and quantitative evaluations, achieving high-fidelity, consistent novel view synthesis at desired camera angles across a wide variety of scenes while maintaining accurate, natural detail representation and image clarity across various viewpoints. We also support our method with a comprehensive analysis of 2D image generation models and the 3D space, providing a solid foundation and rationale for our solution. Furthermore, the proposed framework contributes to scenario generation and ecological visualization by enabling controllable, multi-view synthesis of natural and urban environments from limited imagery. This capability can support climate impact simulations and environmental narrative synthesis, aligning with recent advances in generative AI and foundation models for scientific and ecological applications.
Figure 1: Our model is capable of generating high quality camera-controlled images at specific azimuth and elevation angles for a variety of complex scenes, all without requiring extra 3D datasets or extensive training. The image in the bottom right corner showcases the output from the 3D-based baseline, Zero123++ (Shi et al. 2023a), created from a designated angle.
@@
Introduction
Novel-view synthesis plays a pivotal role in numerous real-world applications, including 3D environments, augmented reality (AR), virtual reality (VR), and autonomous driving. Recent advancements in diffusion-based methods, such as Zero-1-to-3 (Zero123) (Liu et al. 2023a) and Zero123++ (Shi et al. 2023a), along with NeRFs (Tancik et al. 2022; Zhou and Tulsiani 2023; Deng et al. 2023) Gaussian Splatting (Li et al. 2024; Zhu et al. 2023), and so on (Sargent et al. 2023; Tung et al. 2025; Van Hoorick et al. 2025; Burgess, Wang, and Yeung 2023; Wiles et al. 2020; Shen et al. 2021; Tucker and Snavely 2020). They have significantly propelled the field forward. Some techniques enable the specification of camera angles and the sampling of novel-view images from precise viewpoints. However, diffusion-based methods remain largely object-centric and may struggle to generalize to complex scenes with intricate backgrounds. They also require extensive 3D object datasets for training. In contrast, NeRF and Gaussian Splatting methods can handle complex scenes but depend on multi-view information to construct 3D models. Therefore, achieving novel-view synthesis from a single image in a data-efficient manner, without relying on additional 3D, multi-view, or depth information, is highly advantageous.
On the other hand, 3D-free methods such as DreamBooth and other recent models (Kothandaraman et al. 2023b, a; Ruiz et al. 2023) aim to intelligently extract the rich 3D knowledge embedded in text-to-2D image diffusion models, such as stable diffusion (Rombach et al. 2022; Podell et al. 2023), to generate text-controlled views from complex, real-world input images without needing additional multi-view or 3D information for fine-tuning or inference. Among these, HawkI stands out as the current best 3D-free model, demonstrating superior capability in utilizing embedded 3D knowledge for high-quality, text-controlled image synthesis. However, despite its excellence, HawkI, like other 3D-free methods, lacks the ability to precisely control camera angles when generating novel-view images. Ideally, we aim for both data-efficient novel view synthesis and camera controllability, which is the primary focus of this paper.
@@ Introduction
Using these insights, along with the established knowledge that 3D-based methods such as HawkI enable precise camera control and 3D-free methods like Zero123++ offer generalizability and data efficiency, we propose a simple approach for novel-view synthesis that generates camera-controlled novel views at specified azimuth and elevation angles from a single input image, without requiring 3D datasets or extensive training. Our method utilizes information from off-the-shelf pretrained model, specifically using Zero123++ (Shi et al. 2023a), a plug-and-play model, in conjunction with the pretrained stable diffusion model. The process employs a 3D-free HawkI-style optimization procedure during inference to achieve the desired outcomes, utilizing information from 3D-based methods as pseudo or weak guidance images. To improve viewpoint consistency—an area where the CLIP model lacks information—we introduce a regularization loss term. This term promotes alignment between the target angle embedding (which captures elevation and azimuth data) and the optimized embedding. By integrating 3D angular information within the CLIP space through the 3D prior image, we reinforce the specified camera viewpoint in the generated images. We validate our approach through extensive qualitative and quantitative comparisons across various metrics that assess text consistency and fidelity w.r.t. input image.
@@
In summary, the contributions of this paper are as follows: (1) We present a novel approach for novel view synthesis that allows for precise camera control, especially effective for complex images with multiple objects and detailed backgrounds. Our method harnesses insights from pretrained 3D models within a 3D-free framework, removing the necessity for additional multi-view or 3D data during both training and inference, effectively combining the advantages of both approaches. (2) We provide an analysis of the CLIP model’s understanding of 3D space and the role of guidance images in 3D-free methods. This analysis supports our solution by highlighting the importance of using priors from pretrained 3D models to enhance viewpoint information in 3D space, utilizing the 3D prior image as a guiding factor for our task. (3) We present comprehensive qualitative and quantitative results on various synthetic and real images, demonstrating significant improvements over baseline 3D-based and 3D-free methods in terms of text consistency and fidelity relative to the input images. Our model’s results maintains consistent, accurate, natural detail representation and image clarity across various viewpoints. Also, our model outperforms the lowest-performing model by 0.1712 in LPIPS (HawkI-Syn $(-20^{\circ},210^{\circ})$ in Table 3), which is 5.2 times larger than the 0.033 gap of Zero123++ in comparison to the lowest-performing model (Table 1 in (Shi et al. 2023a)).
Prior Work on 3D and 3D-free Approaches for NVS
@@
Recent research has increasingly focused on novel view synthesis using diffusion models (Chen et al. 2021; Mildenhall et al. 2021; Shi, Li, and Yu 2021; Gu et al. 2023). Approaches for 3D generation (Chen et al. 2024; Lin et al. 2023; Poole et al. 2022; Raj et al. 2023; Xu et al. 2023; Gao et al. 2024; Park et al. 2017) often rely on text for reconstruction and require substantial multi-view and 3D data (Shi et al. 2023c; Wang and Shi 2023; Yang et al. 2024; Liu et al. 2023b; Höllein et al. 2024; Jain, Tancik, and Abbeel 2021; Liu et al. 2024; Shi et al. 2023b, c) for supervised learning. For instance, Zero123 (Liu et al. 2023a), Magic123 (Qian et al. 2023), and Zero123++ (Shi et al. 2023a) utilize a pre-trained stable diffusion model (Rombach et al. 2022) combined with 3D data corresponding to 800k objects to learn various camera viewpoints. In other words, they are extremely data hungry, meaning that they need extensive multi-view and 3D data to train on. Additionally, most state-of-the-art view synthesis algorithms are largely object-centric, and may not work well on complex scenes containing multiple objects or background information. This is due to the domain gap between the 3D objects data that they are typically trained on, and the inferencing image.
@@
On the other hand, Free3D (Zheng and Vedaldi 2024) introduces an efficient method for synthesizing accurate 360-degree views from a single image without 3D representations. By incorporating the Ray Conditioning Normalization (RCN) layer into 2D image generators, it encodes the target view’s pose and enhances view consistency with lightweight multi-view attention layers and noise sharing. However, it still requires training on large-scale 3D datasets like Objaverse and multi-view information, and it cannot include background transformations. DreamFusion (Poole et al. 2022) presents a Text-to-3D method using a NeRF and a Diffusion Model-based Text-to-2D model. It introduces a probability density distillation loss, allowing the 2D Diffusion Model to optimize image generation without needing 3D data or model modifications. DreamFusion’s key contributions are creating Text-to-3D models without 3D dataset training and utilizing a Diffusion Model. However, it cannot transform images with backgrounds or introduce elevation changes in camera-controlled images. Aerial Diffusion and HawkI (Kothandaraman et al. 2023a, b) synthesize high-quality aerial view images using text and a single input image without 3D or multi-view information. They employ a pretrained text-to-2D Stable Diffusion model, achieving a balance between aerial view consistency and input image fidelity through test-time optimization and mutual information-based inference. However, Aerial Diffusion doesn’t extend well to complex scenes and has artifacts in the generated results, and HawkI struggles with controlling camera angles, detailed feature generation, and maintaining view consistency.
Understanding 2D models and the 3D space
@@
In this section, we use 3D-free stable diffusion based view synthesis method, HawkI (Kothandaraman et al. 2023b). HawkI employs classical computer vision techniques to generate aerial view images from ground view images through a homography transformation, which acts as the guidance image for the diffusion model. We used the HawkI default setting for experiments, unless stated otherwise.
How well do CLIP models understand the 3D-space?
@@
The main reason stable-diffusion-based 2D models using 3D-free approaches struggle with camera control is their limited understanding of 3D space. While they can grasp concepts like top, bottom, and side, they lack precise camera information, such as “30 degrees to the right.” A key factor in this limitation is the CLIP model, which serves as the vision-language backbone for models like Stable Diffusion. In this section, we analyze how effectively CLIP models comprehend 3D space by examining the 3D-free stable diffusion based view synthesis method, HawkI (Kothandaraman et al. 2023b). Our hypothesis is that HawkI’s capability to execute viewpoint transformations without relying on 3D data is dependent on this homography image.
@@
Importance of guidance image in 3D-free methods
@@
We conduct experiments using the images generated using the pretrained Zero123++ (Shi et al. 2023a) model for guidance. In our experiments, the target text specified the desired transformation angles, but the guidance images had different angles. i.e. the guidance images introduced incorrect viewpoints. When generating an image at a $(30^{\circ},30^{\circ})$ angle, the model followed the guidance image’s suggestion, regardless of the text input. This emphasizes that the model benefits from the information in the 3D-prior model’s guidance image a lot. Our experiment highlights the importance of accurate guidance images for camera control. Figure 4 illustrates these findings.
@@
3D-Free Meets 3D Priors: An approach for 3D-data efficient NVS
According to the analysis, we present a method (Figure 2) for data-efficient text and camera-controlled novel-view synthesis from a single input image ($I_{input}$) and its text description ($t_{input}$) (e.g., “An ancient Egyptian pyramid in the desert,” obtained using the BLIP-2 (Li et al. 2023) model). Our model eliminates the need for training data, 3D data, or multi-view data. Instead, it utilizes a pretrained text-to-2D image stable diffusion model as a strong prior, along with pretrained novel-view synthesis (NVS) models, e.g. Zero123++, for guidance. Our method combines information from the pretrained NVS model and performs a rapid inference-time optimization and inference routine to generate novel-view images of any given in-the-wild complex input scene at specified elevation ($\alpha_{elev}$) and azimuth angles ($\alpha_{azi}$). Elevation ($\alpha_{elev}$) refers to the vertical angle relative to the object, measured in degrees, and is defined based on the orientation of the input image. Similarly, azimuth angles ($\alpha_{azi}$) refer to the horizontal angle around the object, also relative to the input image. We next describe our method in detail.
Inference-time Optimization
We employ a pretrained NVS model $G$ to obtain a weak prediction, $I_{view}$, of $I_{input}$ at ($\alpha_{elev}$, $\alpha_{azi}$). This prediction is represented as $I_{view}=G(I_{input},\alpha_{elev},\alpha_{azi})$. Although $I_{view}$ is not a fully accurate depiction of the desired target, it provides weak or pseudo guidance for the model regarding the content and direction of the desired viewpoint transformation. Subsequently, we utilize the pretrained text-to-image stable diffusion model to perform inference-time optimization (Kothandaraman et al. 2023b).
Across all four steps, the reconstruction loss $L$ is used to guide the optimization process, ensuring accurate reconstruction of $I_{input}$ and $I_{view}$. In Step 4, the addition of the regularization loss reinforces viewpoint consistency by aligning $e_{view}$ with $e_{target}$, thereby improving camera-controlled image generation.
Step 1: Text Embedding Optimization on $I_{input}$
Initially, we enhance the CLIP embedding for $t_{input}$ with $I_{input}$ to derive $e_{optim}$ (optimized CLIP text-image embedding from $e_{input}$, which is the CLIP test embedding for $t_{input}$). This embedding is optimized to most accurately reconstruct $I_{input}$. $f$ represents the diffusion model function that maps the input latent $x_{t}$, timestep $t$, and the optimized embedding $e_{optim}$. The reconstruction is achieved by minimizing the denoising diffusion loss function $L$ (Ho, Jain, and Abbeel 2020), using the frozen diffusion model UNet:
$\it\min_{e_{optim}}\sum_{t=T}^{0}L(f(x_{t},t,e_{optim};\theta),I_{input})$} (1)
@@
Step 2: Fine-tuning UNet on $I_{input}$
@@
Step 3: Text Embedding Optimization on $I_{view}$
@@
$\it\min_{e_{view}}\sum_{t=T}^{0}L(f(x_{t},t,e_{view};\theta),I_{view})$} (3)
Step 4: Fine-tuning UNet on $I_{view}$ with Regularization Loss
@@
$\it\min_{\theta_{LoRA}}\sum_{t=T}^{0}\big(L(f(x_{t},t,e_{view};\theta),I_{view})+L_{reg}\big)$} (4)
Viewpoint Regularization
@@
We camera control results by adding a regularization term between the text embedding that includes elevation and azimuth information ($e_{target}$) and the optimized text embedding ($e_{view}$) in addition to the pretrained guidance model. In addition to enriching the viewpoint knowledge of the 3D space, applying this loss is also essential to address the limitations of the 3D-prior guidance model, as the 3D-prior model does not perform very well in complex scenes. Specifically, by applying a regularization loss between the text embedding that contains angle information and the optimized text embedding, our hypothesis is that we can improve camera control results by building a model that references the guidance image as supplementary information rather than relying solely on it. Hence, to improve viewpoint consistency, an additional regularization loss term is added to the reconstruction loss. This term introduces a constraint between the angle embedding $e_{target}$ (representing elevation and azimuth information) and the optimized embedding $e_{view}$ during this refinement. The regularization term, calculated as $L_{reg}=\|e_{view}-e_{target}\|^{2}$ encourages alignment between $e_{view}$ and the intended angle information in $e_{target}$, reinforcing the target viewpoint in the generated results. Thus, the predicted image from the 3D-based NVS method, Zero123++, serves as weak or pseudo guidance. The optimization strategy conditions the embedding space with knowledge related to the input image and its view variants using the guidance image prior which facilitates view transformation and provides a direction for viewpoint transformation.
Inference To generate the camera-controlled image with designated elevation ($\alpha_{elev}$) and azimuth angles ($\alpha_{azi}$), we use the target text description $t_{target}$, which varies according to the corresponding $\alpha_{elev}$ and $\alpha_{azi}$. For instance, if $\alpha_{elev}=30^{\circ}$ and $\alpha_{azi}=30^{\circ}$, $t_{target}$ can be formatted as ”View from an elevated angle of +30 degrees and an azimuth angle of +30 degrees” + $t_{input}$ (e.g., ”View from an elevated angle of +30 degrees and an azimuth angle of +30 degrees, An ancient Egyptian pyramid in the desert.”). Next, we use the finetuned diffusion model to generate the target image using $t_{target}$, along with mutual information guidance (Kothandaraman et al. 2023b), which enforces similarity between the contents of the generated and input images.
Experiments and Results
@@
We utilize the HawkI-Syn (Kothandaraman et al. 2023b) and HawkI-Real (Kothandaraman et al. 2023b) datasets that feature complex scenes with multiple foreground objects and background. Both datasets provide images and text prompts to the model. While we acknowledge the value of established benchmarks, our chosen datasets, HawkI-Syn and HawkI-Real, cover a broad spectrum of scene types—including indoor/outdoor, human/animal, and synthetic/real-world domains. As shown in prior work (Kothandaraman et al. 2023b), they are specifically designed for evaluating viewpoint generalization in diverse, unbounded settings, making them suitable and representative for our task.
@@
We compare our method with state-of-the-art view synthesis methods: Zero123++ (Shi et al. 2023a) and Stable Zero123 for 3D-based methods and HawkI (Kothandaraman et al. 2023b) for 3D-free method.
@@
We employ the stable diffusion 2.1 model as the backbone for all our experiments and results. To generate the pseudo guidance images for different viewpoints, we use the pretrained Zero123++ (Shi et al. 2023a) model. All images except those in HawkI-Real dataset are used at a resolution of $512\times 512$. For $I_{input}$, we train the text embedding for 1,000 iterations with the learning rate of 1e31e-3 and the diffusion model for 500 iterations with the learning rate of 2e42e-4. Training the text embedding for 1,000 iterations guarantees that the text embedding $e_{optim}$ is not too close to the $e_{input}$, avoiding bias towards $I_{input}$. Likewise, it is not too distant from $e_{input}$, allowing the text embedding space to capture the characteristics of $I_{input}$. Regarding $I_{view}$, we trained the text embedding for 500 iterations and the diffusion UNet for 250 iterations. We aim for $e_{view}$ to be near $e_{optim}$ and limit the diffusion model training to 250 iterations to prevent overfitting to $I_{view}$. The purpose of $I_{view}$ is to introduce variability and provide pseudo supervision rather than accurately approximating the camera control. We set the mutual information guidance hyperparameter to 1e61e-6 and conduct inference over 50 steps.
Qualitative analysis
@@
$\{(\alpha_{elev},\alpha_{azi})\mid(30^{\circ},30^{\circ}),(-20^{\circ},210^{\circ}),$
$(30^{\circ},270^{\circ}),(-20^{\circ},330^{\circ})\}.$